OpenStudy (anonymous):

Which is 10–4 written in standard form? A. 0.00001 B. 0.0001 C. 1000 D. 10,000

OpenStudy (anonymous):



OpenStudy (anonymous):

10^-4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What is normally in the fourth's place?

OpenStudy (anonymous):

ones, tens, hundreds, thousands?

OpenStudy (anonymous):

thousands

OpenStudy (ranga):

\[\Large 10^{-4} = \frac{ 1 }{ 10^4 } = \frac{ 1 }{ 10000 } = ?\]

OpenStudy (anonymous):

right, so you want to create a decimal with 10 or 1 in the thousands place

OpenStudy (anonymous):

.0001

OpenStudy (anonymous):

b

OpenStudy (anonymous):

Right, do you understand why?

OpenStudy (anonymous):

yes i divided 1 and 10000

OpenStudy (anonymous):

Right, good job! Good luck!